Use 2 methods to show triangle with vertices A (-6, 5), B (2, 1), C (-2, -6) is a right triangle.
*Analytic Geometry Question*

Answer by Alan3354(69443)

Use 2 methods to show triangle with vertices A (-6, 5), B (2, 1), C (-2, -6) is a right triangle.
-------------
Method 1:
Find the slope of AB, BC and AC.
If any pair is perpendicular, it's a right triangle.
----
Method 2:
Find the length of the 3 sides.
If c^2 = a^2 + b^2 it's a right triangle (c = longest side).